OpenStudy (anonymous):

When solving negative one over eight (x + 35) = −7, what is the correct sequence of operations? Multiply each side by negative one over eight , add 35 to each side Multiply each side by negative one over eight , subtract 35 from each side Multiply each side by −8, subtract 35 from each side Multiply each side by −8, add 35 to each side

OpenStudy (michele_laino):

If I multiply both sides, by \(−8\) I get this: \[\left( { - 8} \right) \cdot \left( { - \frac{1}{8}} \right)\left( {x + 35} \right) = \left( { - 8} \right) \cdot \left( { - 7} \right)\]

OpenStudy (michele_laino):

please simplify

OpenStudy (anonymous):

so would it be A Multiply each side by negative one over eight , add 35 to each side

OpenStudy (michele_laino):

please nothe that I have multiplied by \(-8\) not by \(-1/8\)

OpenStudy (michele_laino):

note*

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

no i mean c

OpenStudy (michele_laino):

correct! since after that multiplication, I can write this: \[x + 35 = 56\] so, If I subtract \(35\) I get this: \[x + 35 - 35 = 56 - 35\] or: \(x=21\)

OpenStudy (anonymous):

thx you helped me big time

OpenStudy (michele_laino):

:)